EXN Status Page
All Systems Operational
EXtraNet Portal   ? Operational
Shared Hosting   ? Operational
Layer 2 Services   ? Operational
IP Transit Services   ? Operational
Leased Lines   Operational
Virtual Data Centre (vDC) - Global Switch 2   ? Operational
Virtual Data Centre (vDC) - Virtus Enfield   ? Operational
Datacentres Operational
Global Switch 2   Operational
Reading   Operational
Telecity Meridian Gate   Operational
Virtus Enfield   Operational
Virtus Hayes   Operational
Volta GSS   Operational
Telehouse North   Operational
Telehouse East   Operational
Virtus Slough   Operational
Operational
Degraded Performance
Partial Outage
Major Outage
Maintenance
Scheduled Maintenance
EXN Christmas 2017 Change Freeze Dec 15, 13:00 - Jan 4, 09:00 GMT
EXN will be implementing an change freeze from Friday 15/12/2017 at 13:00 GMT until Thursday 04/01/2018 at 09:00 GMT. This freeze is implemented every year during the busy Christmas trading period to ensure that no changes are made to our network or infrastructure which could affect production services.

The freeze includes all changes to the network and non-essential maintenance or infrastructure changes. EXN won't be undertaking any planned maintenance during this time, unless we are responding to a fault.

Fault management and emergency tickets will not be affected by this change.

If you have any further questions regarding this freeze, please get in touch via the usual channels.
Posted on Oct 27, 11:23 BST
Past Incidents
Nov 22, 2017

No incidents reported today.

Nov 21, 2017

No incidents reported.

Nov 20, 2017

No incidents reported.

Nov 19, 2017
Completed - The scheduled maintenance has been completed.
Nov 19, 00:01 GMT
In progress - Scheduled maintenance is currently in progress. We will provide updates as necessary.
Nov 17, 00:00 GMT
Scheduled - VIRTUS LONDON1 (Enfield) have informed us of a generator delivery which will take place between Saturday 18th November and Sunday 19th November 2017.

A crane will be used to lift the generator into position, this may take up some of the customer car parking spaces, which will limit the parking spaces available on site. However, Virtus will try to accommodate you with a parking space during the delivery times listed above. We also advise all customers to take extra care when moving around the car park when the delivery is taking place.
Nov 16, 10:54 GMT
Nov 18, 2017
Completed - We have been informed by the data centre that this maintenance has been postponed. Another maintenance window will be announced in due course.
Nov 18, 19:08 GMT
Update - Please be advised that the transfer back planned for this Saturday has been postponed for 2 weeks until Saturday 18th November at 23:00.
Nov 3, 09:52 GMT
Scheduled - On Saturday 4th November Virtus are completing essential UPS maintenance. During this time the "B" feed of this suite will be switched back to raw mains causing a very short outage on the "B" feed.

If you have any queries, please do not hesitate to contact us quoting the ticket reference number T20170915.0050.
Oct 30, 10:40 GMT
Resolved - EXN have had the following update from Global Switch:

At 10:58 on 18 November 2017, Global Switch London East Level 10 Package Sub Station ‘B’ (PSS10B) Main Incoming Low Voltage (LV) Air Circuit Breaker (ACB) tripped causing a Loss of the ‘B’ feed to all suites on Level 10.

The Power Loss was picked up by our Power Management System (PMS) in the LONE Control Room. The Global Switch Authorised Person (AP) confirmed that PSS 10 A was healthy and that PSS 10 B Incoming LV ACB had opened. All out going circuit breakers on PSS 10 B were opened and the Main Incoming LV ACB was reclosed. Once Power was restored to the 10 B LV Switchboard all out going circuits were systematically reclosed. A full sweep of all suites then followed to check all DFU’s, PDU’s and STS’s. The Incident concluded at approximately 11:35.

We are expecting a full RCA from Global Switch in the coming days and this will be shared with customers on an individual basis.
Nov 18, 15:02 GMT
Update - Further internal investigations indicate that this issue was isolated to the "B" feed of rows "B" and "C" in suite J03 (10th floor) in GS2. Clients in the "D" row and in suite G04 (7th floor) were not affected. Updates to follow.
Nov 18, 11:47 GMT
Update - Both power feeds are now operational, we are awaiting updates from Global Switch.
Nov 18, 11:29 GMT
Update - A major incident has been raised with the data centre who are investigating the issue. At this stage we don't have any details of the nature of the problem but we expect an update within the next 30 minutes.
Nov 18, 11:21 GMT
Update - At this time this outages appears to be on a single feed. All network and VDC services remain up.
Nov 18, 11:04 GMT
Investigating - We are currently investigating a power outage within Global Switch 2. Further updates to follow.
Nov 18, 11:01 GMT
Nov 16, 2017

No incidents reported.

Nov 15, 2017

No incidents reported.

Nov 14, 2017

No incidents reported.

Nov 13, 2017

No incidents reported.

Nov 12, 2017

No incidents reported.

Nov 11, 2017

No incidents reported.

Nov 10, 2017

No incidents reported.

Nov 9, 2017

No incidents reported.

Nov 8, 2017

No incidents reported.